Solve using the quadratic formula:
x2 – 3x – 23 = 5

Answer by checkley77(12844)

x^2–3x–23=5
x^2-3x-23-5=0
x^2-3x-28=0
(x-7)(x+4)=0
x-7=0
x=7 ans.
x+4=0
x=-4 ans.